# Early Failure Detection in Autonomous Surgical Soft-Tissue Manipulation via Uncertainty Quantification

### v1

@@ Abstract
Autonomous surgical robots are a promising solution to the increasing demand for surgery amid a shortage of surgeons. Recent work has proposed learning-based approaches for the autonomous manipulation of soft tissue. However, due to variability in tissue geometries and stiffnesses, these methods do not always perform optimally, especially in out-of-distribution settings. We propose, develop, and test the first application of uncertainty quantification to learned surgical soft-tissue manipulation policies as an early identification system for task failures. We analyze two different methods of uncertainty quantification, deep ensembles and Monte Carlo dropout, and find that deep ensembles provide a stronger signal of future task success or failure. We validate our approach using the physical daVinci Research Kit (dVRK) surgical robot to perform physical soft-tissue manipulation. We show that we are able to successfully detect task failure and request human intervention when necessary while still enabling autonomous manipulation when possible. Our learned tissue manipulation policy with uncertainty-based early failure detection achieves a zero-shot sim2real performance improvement of 47.5% over the prior state of the art in learned soft-tissue manipulation. We also show that our method generalizes well to new types of tissue as well as to a bimanual soft-tissue manipulation task.
@@ I Introduction
Autonomous surgical robots have the potential to help solve the growing disparity between the population’s need for surgery and the number of available surgeons [3, 4]. However, surgical robot learning and automation is particularly challenging due to the nuanced and risk-sensitive nature of the tasks, the partially observable and deformable nature of the environment, and the scarcity of available data. In particular, because autonomous surgical system failures can be detrimental to patient health, it is crucial that these autonomous systems take into account uncertainty so that the system can safely cede control to the surgeon before causing any harm. Therefore, we aim to develop an autonomous surgical soft-tissue manipulation system that is capable of reasoning over predictive uncertainty to detect task failure early in the robot’s operation (see Fig. 1). In doing so, our goal is to mitigate the risk of system failures while still offloading the work of soft-tissue manipulation to the robot, whenever it is safe to do so.
@@
Our main goal is to enable uncertainty-based task failure prediction in the domain of soft-tissue manipulation. To this end, we build on recent advances in robot learning for deformable object manipulation. In particular, we focus on the recently proposed, state-of-the-art DeformerNet framework [2] which uses large-scale self-supervised training in simulation to learn a closed-loop policy for manipulating deformable objects to desired goal geometries. While DeformerNet has been shown to perform well on in-distribution tissue, inter-human variability can lead to erroneous outputs, which can pose a safety risk. We propose to address this issue through quantifying the predictive uncertainty of DeformerNet, eliminating potentially incorrect and harmful manipulations of the tissue.
@@
One of the major obstacles when applying robot learning to surgical domains is the inability to do online learning in the real world. While we can collect demonstrations offline [5, 6, 7, 8] and can train policies in simulation [9, 10, 11, 12], offline imitation learning leads to compounding errors [13] and the sim2real gap is especially difficult to overcome for the types of deformable manipulation that are common in surgery [14]. Enabling a surgical robot to preemptively detect task failures addresses these challenges by enabling a human to correct the robot during real-world execution, thus preventing failures while also enabling the robot to potentially use human interventions to continually improve its policy and world model. While there has been some prior work on interactive imitation learning and human interventions [15, 16, 17, 18, 19, 20], to the best of our knowledge, we are the first to study robot-requested interventions on the kinds of complex deformable object manipulation tasks common in surgical robotics.
@@
Our main contributions are as follows: (1) We develop and study the first real-world surgical robotic system capable of uncertainty quantification and early task failure detection during soft-tissue manipulation. (2) We extend and analyze two potential methods for quantifying predictive uncertainty during learned soft-tissue manipulation: deep ensembles and Monte Carlo dropout. We provide empirical evidence that ensembles provide more accurate uncertainty estimates and that using the slope of the ensemble variance rather than the raw variance more accurately discriminates between successful and unsuccessful tissue manipulation actions. (3) Across 40 real-world tissue manipulation trials, we show that our approach improves zero-shot sim2real performance of more than 47% when compared with DeformerNet by itself, the prior state-of-the-art approach for deformable tissue manipulation. Concretely, our method achieves a 95% system success rate compared with DeformerNet achieving 47.5% system success rate without our method. Further, our learned thresholds for failure detection successfully generalize to new types of biological tissue and to tasks unseen during training and calibration.
@@ II-B Uncertainty Quantification
Monte Carlo dropout is another technique utilized for uncertainty estimation in deep neural networks [35, 36]. Traditional dropout regularization, initially introduced to mitigate overfitting during training, randomly deactivates units in a neural network, effectively creating an ensemble of different network architectures. By extending dropout to the inference phase and making multiple inference passes through the network with different dropout instantiations, Monte Carlo dropout can be used to generate a distribution of outputs. Unlike ensembles, which require training multiple independent models, Monte Carlo dropout leverages the same trained model for inference, eliminating the need for repeated training [37].
By analyzing this distribution from an ensemble or Monte Carlo dropout, uncertainty metrics such as entropy or variance across predictions can be computed, providing insights into the model’s confidence and reliability [38, 39].
@@ III Problem Definition
The meta policy $\pi_{\rm meta}$ chooses whether to execute the learned robot policy or to hand over control to a human supervisor and can thus be represented as a binary classifier, $\pi_{\rm meta}:\mathcal{O}\rightarrow\{0,1\}$, where $\mathcal{O}$ is the observation space. A true positive (TP) is when the robot correctly requests an intervention and would have failed if it had operated autonomously, a false positive (FP) is when the robot wrongly asks for an intervention when it would have succeeded autonomously, a false negative (FN) is when the robot does not ask for an intervention and fails at the task, and a true negative (TN) is when the robot does not request an intervention and succeeds at the task autonomously. There are many valid ways to tune the thresholds of the meta controller to appropriately balance true positives, false positives, true negatives, and false negatives. We describe one possible approach in the following paragraphs.
@@ IV-A DeformerNet
For this work, we modify and endow the DeformerNet framework [2] with uncertainty awareness and the ability to preemptively detect task failures conditioned on that uncertainty. DeformerNet is a deep learning based shape-servoing model for manipulating deformable objects. The model takes as input the current geometry $\mathrm{P}_{c}$ and goal geometry $\mathrm{P}_{g}$ of the deformable object as partial-view point clouds along with the manipulation point $\boldsymbol{m}$ on the object. The model then predicts an action $\hat{\mathrm{A}}$ as a homogeneous transformation matrix for the robot’s end-effector, $\hat{\mathrm{A}}=\begin{bmatrix}\hat{\mathrm{R}}&\hat{\boldsymbol{p}}\\ \boldsymbol{0}&1\end{bmatrix}$, where $\hat{\mathrm{R}}$ and $\hat{\boldsymbol{p}}$ are the predicted change in end-effector orientation and position respectively. DeformerNet is a closed loop control method that reasons over features of the deformable object’s geometry and the goal shape’s geometry leveraging encoders based on PointConv [55]. DeformerNet is trained in a self-supervised fashion using large-scale simulated data. In simulation, given a manipulation point $\boldsymbol{m}$ on a deformable object with current geometry $\mathrm{P}_{c}$, a random homogeneous transformation matrix A is applied to the robot’s end effector. The resulting geometry of the deformable object is then sensed and the new geometry is used as a self-supervised goal, $\mathrm{P}_{g}$. DeformerNet is trained to predict the homogeneous transformation that led to $\mathrm{P}_{g}$ over a large self-supervised data set consisting of instances of the following form: $(\mathrm{P}_{c},\mathrm{P}_{g},\boldsymbol{m},\mathrm{A})$. DeformerNet has been shown to be capable of automating complex deformable object manipulation tasks, including surgical subtasks, with zero-shot sim2real transfer [2].
@@ IV-B Uncertainty quantification
We compute the ensemble’s predictive variance for both position and rotation individually. Positional variance is measured as the average squared distance to the deep ensemble’s positional prediction ($\hat{\boldsymbol{p}}$ in Equation (2)) for each component network:
@@
We define the predictive uncertainty of DeformerNet at time-step $t$ of the trajectory as a vector containing the change in both positional and rotational variances from equations (4) and (5).
@@ IV-C Monte Carlo Dropout
We also investigate uncertainty quantification using Monte Carlo dropout. Using stochastic dropout at inference time, each forward pass through the network for a given input results in a variable output. This enables efficient sampling of a distribution of potential changes in end-effector pose. We sample a distribution of outputs using Monte Carlo dropout and use equations (2) – (6) to get the model predictions and variances.
@@ V Physical Experimental Setup
We train a deep ensemble DeformerNet consisting of five models. We use the source code and training and testing datasets provided by Thach et al. [2], consisting of 11,566 training and 1,285 test examples of manipulations of a deformable box object. Following prior work [31, 34, 32], each ensemble component model was trained using the same training dataset with different random weight initializations. Using the same training and testing datasets as described for the deep ensemble, we train a Monte Carlo dropout version of the DeformerNet model with dropout added to the last layer at varying levels of dropout percentage: 25%, 50%, and 75%. We collect 100 samples from the model to construct the predictive distributions.
@@
To segment the point clouds of the tissue geometry, we leverage the state-of-the-art Segment Anything Model (SAM) [1]. SAM takes as input an RGB image along with a set of query points on the desired segmented object and outputs a segmentation mask for the desired object. We then map the segmentation mask to the depth image of the scene. Using the segmented depth image with the camera’s intrinsic matrix enables us to generate a point cloud on the surface of the desired geometry. The perception pipeline is visualized in Fig. 1.
@@ VI-B Uncertainty Quantification
We collected 30 trials containing both in-distribution and out-of-distribution cases of DeformerNet on ex vivo chicken muscle tissue and measured the predictive variances across the trajectories using Equations 4 and 5. Figure 2 shows the distributions of these raw variance values throughout the successful and unsuccessful trials. We find that raw variance values are not good indicators of task success, as there is no clear distinction between the successful and unsuccessful trials.
@@
To detect task failures, we set a threshold on the slope of the positional and rotational variances at time-step 1 in the trajectories ($\boldsymbol{u}(1)$ from Eq. 6) by calibrating to desired false positive and negative rates on the prior 30 trials where false positives are detecting task failure in successful cases, and false negatives are failing to detect task failure when necessary. False positive and negative rates for different threshold values on 20 test trials can be seen in Fig. 3.
@@
To test the generalization of our method to new tissue geometries and densities, we ran 20 trials on ex vivo bovine tissue with no re-calibration of the chosen thresholds. The results for the bovine tissue are summarized in Table I. We successfully classify task completion in 18 and 17 of the trials for positional and rotational intervention respectively while only failing to detect a task failure in 1 of the positional intervention trials. When using both positional and rotational thresholds, we successfully classify task completion in 17 trials and never fail to detect a task failure.
@@ VII Conclusion
Robot learning for the autonomous completion of surgical tasks is a particularly challenging domain due to data scarcity and high levels of risk. We introduce a novel framework for the safe implementation of an autonomous soft-tissue manipulation system. Leveraging the state-of-the-art deformable object manipulator DeformerNet, we show how a deep ensemble can be used to detect task failure in advance. We apply our method on the dVRK surgical robot system for a soft-tissue manipulation task using both ex vivo chicken and bovine tissue and demonstrate with high accuracy the ability to request interventions in cases that would have otherwise failed.
While this work focuses on early task failure detection, in the future we plan to study bi-direction hand-offs by having the robot continually estimate model uncertainty during the interventions, allowing the robot to indicate to the human a readiness to resume autonomous operation. We also plan to study when humans feel confident ceding control back to the robot and how to visualize model uncertainty to enhance interpretability and trust. Our work assumes that the distributions of outputs are unimodal. It is possible, however, that the true distributions are multimodal. We plan to relax this assumption in the future to dynamically fit a distribution to the predictions that can account for this potential multimodality.

### v2

@@ Abstract
Autonomous surgical robots are a promising solution to the increasing demand for surgery amid a shortage of surgeons. Recent work has proposed learning-based approaches for the autonomous manipulation of soft tissue. However, due to variability in tissue geometries and stiffnesses, these methods do not always perform optimally, especially in out-of-distribution settings. We propose, develop, and test the first application of uncertainty quantification to learned surgical soft-tissue manipulation policies as an early identification system for task failures. We analyze two different methods of uncertainty quantification, deep ensembles and Monte Carlo dropout, and find that deep ensembles provide a stronger signal of future task success or failure. We validate our approach using the physical daVinci Research Kit (dVRK) surgical robot to perform physical soft-tissue manipulation. We show that we are able to successfully detect out-of-distribution states leading to task failure and request human intervention when necessary while still enabling autonomous manipulation when possible. Our learned tissue manipulation policy with uncertainty-based early failure detection achieves a zero-shot sim2real performance improvement of 47.5% over the prior state of the art in learned soft-tissue manipulation. We also show that our method generalizes well to new types of tissue as well as to a bimanual soft-tissue manipulation task.
@@ I Introduction
Autonomous surgical robots have the potential to help solve the growing disparity between the population’s need for surgery and the number of available surgeons [55, 5]. However, surgical robot learning and automation is particularly challenging due to the nuanced and risk-sensitive nature of the tasks, the partially observable and deformable nature of the environment, and the scarcity of available data. In particular, because autonomous surgical system failures can be detrimental to patient health, it is crucial that these autonomous systems take into account uncertainty so that the system can safely cede control to the surgeon before causing any harm. Thus, we aim to develop an autonomous surgical soft-tissue manipulation system that is capable of reasoning over predictive uncertainty to detect task failure early in the robot’s operation (see Fig. 1). In doing so, our goal is to mitigate the risk of system failures while still offloading the work of soft-tissue manipulation to the robot, whenever it is safe to do so.
@@
Our main goal is to enable uncertainty-based task failure prediction in the domain of soft-tissue manipulation. To this end, we build on recent advances in robot learning for deformable object manipulation. In particular, we focus on the recently proposed, state-of-the-art DeformerNet framework [50] which uses large-scale self-supervised training in simulation to learn a closed-loop policy for manipulating deformable objects to desired goal geometries. While DeformerNet has been shown to perform well on in-distribution tissue, inter-human variability can lead to erroneous outputs, which can pose a safety risk. We propose to address this issue through quantifying the predictive uncertainty of DeformerNet.
@@
One of the major obstacles when applying robot learning to surgical domains is the inability to do online learning in the real world. While we can collect demonstrations offline [40, 51, 12, 54] and can train policies in simulation [56, 36, 21, 46], offline imitation learning leads to compounding errors [43] and the sim2real gap is especially difficult to overcome for the types of deformable manipulation that are common in surgery [16]. Enabling a surgical robot to preemptively detect task failures addresses these challenges by enabling a human to correct the robot during real-world execution, thus preventing failures while also enabling the robot to potentially use human interventions to continually improve its policy and world model. While there has been prior work on interactive imitation learning and human interventions [22, 53, 48, 30, 18, 8], to the best of our knowledge, we are the first to study robot-requested interventions on the kinds of complex deformable object manipulation tasks common in surgical robotics.
@@
Our main contributions are as follows: (1) We develop and study the first real-world surgical robotic system capable of uncertainty quantification and early task failure detection during soft-tissue manipulation. (2) We extend and analyze two potential methods for quantifying predictive uncertainty during learned soft-tissue manipulation: deep ensembles and Monte Carlo dropout. We provide empirical evidence that ensembles provide more accurate uncertainty estimates and that using the slope of the ensemble variance rather than the raw variance more accurately discriminates between successful and unsuccessful tissue manipulation actions. (3) Across 40 real-world tissue manipulation trials, we show that our approach improves zero-shot sim2real performance of more than 47% when compared with DeformerNet by itself, the prior state-of-the-art approach for deformable tissue manipulation. Concretely, our method achieves a 95% system success rate compared with DeformerNet achieving 47.5% system success rate without our method. Further, our learned thresholds for failure detection successfully generalize to out-of-distribution tissue and to new tasks unseen during training and calibration.
@@ II-B Uncertainty Quantification
Monte Carlo dropout is another technique utilized for uncertainty estimation in deep neural networks [33, 6]. By extending dropout to the inference phase and making multiple inference passes through the network with different dropout instantiations, Monte Carlo dropout can be used to generate a distribution of outputs. Unlike ensembles, which require training multiple independent models, Monte Carlo dropout leverages the same trained model for inference, eliminating the need for repeated training [29].
By analyzing the distribution from an ensemble or Monte Carlo dropout, uncertainty metrics such as entropy or variance across predictions can be computed, providing insights into the model’s confidence and reliability [45, 31].
@@ III Problem Definition
The meta policy $\pi_{\rm meta}$ chooses whether to execute the learned robot policy or to hand over control to a human supervisor and can thus be represented as a binary classifier, $\pi_{\rm meta}:\mathcal{O}\rightarrow\{0,1\}$, where $\mathcal{O}$ is the observation space. A true positive (TP) is when the robot correctly requests an intervention and would have failed if it had operated autonomously, a false positive (FP) is when the robot wrongly asks for an intervention when it would have succeeded autonomously, a false negative (FN) is when the robot does not ask for an intervention and fails at the task, and a true negative (TN) is when the robot does not request an intervention and succeeds at the task autonomously.
@@ IV-A DeformerNet
For this work, we modify and endow the DeformerNet framework [50] with uncertainty awareness and the ability to preemptively detect task failures conditioned on that uncertainty. DeformerNet is a deep learning based shape-servoing model for manipulating deformable objects. The model takes as input the current geometry $\mathrm{P}_{c}$ and goal geometry $\mathrm{P}_{g}$ of the deformable object as partial-view point clouds along with the manipulation point $\boldsymbol{m}$ on the object. The model then predicts an action $\hat{\mathrm{A}}$ as a homogeneous transformation matrix for the robot’s end-effector, $\hat{\mathrm{A}}=\begin{bmatrix}\hat{\mathrm{R}}&\hat{\boldsymbol{p}}\\ \boldsymbol{0}&1\end{bmatrix}$, where $\hat{\mathrm{R}}$ and $\hat{\boldsymbol{p}}$ are the predicted change in end-effector orientation and position respectively. DeformerNet is a closed loop control method that reasons over the deformable object’s geometry and the goal shape’s geometry leveraging encoders based on PointConv [52]. DeformerNet is trained in a self-supervised fashion using large-scale simulated data. In simulation, given a manipulation point $\boldsymbol{m}$ on a deformable object with current geometry $\mathrm{P}_{c}$, a random homogeneous transformation matrix A is applied to the robot’s end effector. The geometry of the deformable object is then sensed and is used as a self-supervised goal, $\mathrm{P}_{g}$. DeformerNet is trained to predict the homogeneous transformation that led to $\mathrm{P}_{g}$ over a data set consisting of instances of the following form: $(\mathrm{P}_{c},\mathrm{P}_{g},\boldsymbol{m},\mathrm{A})$.
@@ IV-B Uncertainty quantification
We compute the ensemble’s predictive variance for both position and rotation individually. Positional variance is the average squared distance to the deep ensemble’s positional prediction ($\hat{\boldsymbol{p}}$ in Equation (2)) for each component network:
@@
We define the predictive uncertainty of DeformerNet at time-step $t$ as a vector containing the change in both positional and rotational variances from equations (4) and (5).
@@ IV-C Monte Carlo Dropout
We also investigate uncertainty quantification using Monte Carlo dropout. Using stochastic dropout at inference time allows us to sample from a distribution of potential changes in end-effector pose. We then use equations (2) – (6) to get the model predictions and variances.
@@ V Physical Experimental Setup
We train a deep ensemble DeformerNet consisting of five models. We use the source code and training and testing datasets provided by Thach et al. [50], consisting of 11,566 training and 1,285 test examples of manipulations of a deformable box object. Following prior work [27, 25, 13], each ensemble component model was trained using the same training dataset with different random weight initializations. We train a Monte Carlo dropout version of the DeformerNet model with dropout added to the last layer at varying levels of dropout percentage: 25%, 50%, and 75%. We collect 100 samples from the model to construct the predictive distributions.
@@ VI-B Uncertainty Quantification
We collected 30 trials containing both in-distribution and out-of-distribution cases of DeformerNet on ex vivo chicken muscle tissue and measured the predictive variances across the trajectories using Equations 4 and 5. Figure 2 shows the distributions of these raw variance values for successful and unsuccessful trials. We find that raw variance values are not good indicators of task success, as there is no clear distinction between the successful and unsuccessful trials.
@@
To detect task failures, we set a threshold on the slope of the positional and rotational variances at time-step 1 in the trajectories ($\boldsymbol{u}(1)$ from Eq. 6) by calibrating to desired false positive and negative rates on the above 30 trials (false positives are detecting task failure in successful cases, false negatives are failing to detect task failure when necessary). False positive and negative rates for different threshold values on 20 test trials can be seen in Fig. 3.
@@
To test the generalization of our method to new tissue types, we ran 20 trials on ex vivo bovine tissue with no re-calibration of the chosen thresholds. The results for the bovine tissue are summarized in Table I. We successfully classify task completion in 18 and 17 of the trials for positional and rotational intervention respectively while only failing to detect a task failure in 1 of the positional intervention trials. When using both thresholds, we successfully classify task completion in 17 trials and never fail to detect a task failure.
@@ VII Conclusion
Robot learning for the autonomous completion of surgical tasks is a particularly challenging domain due to data scarcity, out-of-distribution observations, and safety concerns. To address this challenge, we introduce a novel framework for the safe implementation of an autonomous soft-tissue manipulation system for uncertainty quantification and early failure detection. We apply our method on the dVRK surgical robot system for a soft-tissue manipulation task using both in- and out-of-distribution ex-vivo tissue types and demonstrate with high accuracy the ability to request interventions in cases that would have otherwise failed. In the future we plan to study bi-directional human-robot hand-offs during soft tissue manipulation as well as study when humans feel confident ceding control back to the robot and how to visualize model uncertainty to enhance interpretability and trust.
Acknowledgments
Research reported in this publication was supported by the Advanced Research Projects Agency for Health (ARPA-H) under Award Number D24AC00415-00. The ARPA-H award provided 100% of the total costs with an award total of up to \$11,935,038. The content is solely the responsibility of the authors and does not necessarily represent the official views of ARPA-H.